TEST ARTICLE

1. First Section

Nothing going on.

2. Second Section

2.1. Subsection eins. 2.2. Subsection zwei. Theorem 2.1. Initial theorem. Proof. The proof includes cases after a lengthy and somewhat pointless discourse that is long enough to wrap around at least to a second line.

Case 1. This is the first case.

Case 2. This is the second case. This then concludes the proof. □

Then in the regular text we trip over more cases. Case 1. Here is a regular text case … Case 2…. and another … Case 3…. and another.

Now let’s see if they work in the statement of a proposition. Lemma 2.2. Silly things happen in one of two ways:

Case 1. first possible pathology; else Case 2. second possible pathology.

Proof. blah blah blah □ Enough is enough.